# Occlusion-Aware Ground Target Tracking by a Dubins Vehicle Using Visibility Volumes

### v1

@@ Abstract
This paper considers the problem of tracking a point of interest (POI) moving along a known trajectory on the ground with an uncrewed aerial vehicle (UAV) modeled as a Dubins vehicle using a line-of-sight (LOS) sensor through an urban environment that may occlude the POI. A visibility volume (VV) encodes a time-varying, three-dimensional representation of the sensing constraints for a particular POI position. A constant-altitude, translating, and radially time-varying circular standoff orbit is then inscribed within the dynamically changing VV centered at the POI position. The time-varying VV is approximated by placing static VVs along the POI’s trajectory using an adaptive metric that restricts the volume change of consecutive visibility volumes to below a specified rate. The time-varying circular standoff orbit is proven to be feasible for a Dubins vehicle and is approximated with a piecewise set of linearly interpolated circular orbits inside the static VVs. A steering controller is derived that drives the UAV to converge to the time-varying standoff orbit. Numerical simulations and a flight test illustrate the proposed approach.
@@ 1 Nomenclature
$(x,y,z)$  = inertial directions easting, northing, and up
@@
$\psi$  = yaw angle about $z$ direction
$\mathbb{S}$  = set of all angles; $[0,2\pi)$
$t$  = time
$k$  = time step
@@
$f(\bm{q},u_{\psi})$  = the dynamics of the UAV
$v$  = forward velocity of the UAV
$u_{\psi}$  = turn-rate control for the UAV
$u_{\psi,{\rm max}}$  = maximum turn rate control for the UAV
$\kappa$  = curvature of a trajectory
$\kappa_{s}$  = signed curvature of a trajectory
$\kappa_{\rm max}$  = maximum curvature of the UAV
$r_{\rm min}$  = minimum turn radius of the UAV; $1/\kappa_{\rm max}$
$h_{\rm UAV}$  = flight altitude of the UAV
$N_{B}$  = number of 2.5-dimensional occluding objects
$B_{i}$  = $i$th 2.5-dimensional object, an extruded polygon
$h_{i}$  = height of $B_{i}$
$B$  = set of 2.5-dimensional objects; $\{B_{0},\ldots,B_{N_{B}-1}\}$
$A_{i}$  = order set of points representing the base of $B_{i}$
@@
$N$  = set of nodes at road intersections
$E$  = set of edges connecting road vertices $E$
@@
$\bm{g}$  = inertial position of the moving point of interest; $[g_{x},g_{y}]^{\rm T}$
$v_{\rm g}$  = magnitude of the velocity for the moving point of interest
$\gamma$  = direction of the point of interest’s velocity
@@
$\bm{p}$  = sampled point along the ground plane $\bm{p}\in\mathbb{R}^{2}$
@@
$d_{\rm max}$  = maximum viewing distance for imaging
$\alpha(\bm{g})$  = visibility volume for target position $\bm{g}$
$\hat{\alpha}(\bm{g})$  = visibility volume approximated as a triangular mesh
$h_{\rm building}$  = maximum height of the tallest building
$h_{\rm feasible}$  = maximum height of the feasible airspace
@@
$\bm{\rho}$  = point in $\mathbb{R}^{3}$
$L(\tau;\bm{g},\bm{\rho})$  = equation of a line from $\bm{g}$ to $\bm{\rho}$ parameterized by $\tau$
$P$  = polygon
@@
$R(t)$  = radius of a morphing visibility orbit
$\theta$  = angle about the point $\bm{g}$ measured from a line parallel to the $x$ axis
$\bm{x}_{\mathcal{I}}$  = inertial trajectory of the UAV in the $xy$ plane $\bm{x}_{\mathcal{I}}\in\mathbb{R}^{2}$
$\bm{e}_{r}$  = radial direction from the point $\bm{g}$ to the UAV at $\bm{q}$ in the $xy$ plane
$\bm{e}_{\theta}$  = direction perpendicular to $\bm{e}_{r}$ in the $xy$ plane
$r$  = UAV’s distance from the POI
$\tilde{r}$  = radial error from the orbit $O(\theta,t)$
$J$  = cost function corresponding to the curvature $\kappa$
$d(\hat{\alpha}(\bm{g}_{i}),\hat{\alpha}(\bm{g}_{j}))$  = measure of the difference between two triangular meshes approximating visibility volumes
[…]
$d_{\rm cutoff}$  = maximum metric between two visibility volumes
$V_{1}$  = a strictly positive scalar function excluding configurations on the orbit $O(\theta,t)$
@@
$\bm{\xi}$  = inertial coordinates for a kinematic particle; $\xi\in\mathbb{R}^{2}$
$\bm{u}$  = a vector field that forces a particle onto an orbit $O(\theta,t)$; $\bm{u}\in\mathbb{R}^{2}$
$\bm{M}$  = a matrix component of $\bm{u}$
@@
$\psi_{d}$  = desired heading angle
@@
$V_{2}$  = a Lyapunov-like function
$\tilde{\psi}$  = angular error from the desired heading $\psi_{d}$
@@
UAV  = uncrewed aerial vehicle
LOS  = line of sight
POI  = point of interest
VV  = visibility volume
VP  = visibility polygon
VO  = visibility orbit
@@ 2 Introduction
Uncrewed aerial vehicles (UAVs) are widely used for airborne sensing tasks that require maintaining a direct line of sight (LOS) to a target point of interest (POI). Applications of tracking a moving POI include cinematography, filming sporting events and parades, search and rescue, and police and military surveillance [1, 2, 3, 4]. For UAV tracking missions where a POI moves along a road through an urban environment, the directional sensor’s view of a POI may be obstructed by buildings, tunnels, bridges, trees, and other structures. This work creates a planning and control method to ensure a UAV maintains an unobstructed LOS to the a POI, moving along a known trajectory, while continuously orbiting it. By orbiting the POI, the UAV can observe it from various angles. Moreover, an encircling motion is required when the UAV cannot hover over the POI (e.g., if the POI’s speed is less than the UAV’s minimum speed). The POI moves over a road network represented as a graph with nodes representing points along roads and edges representing road segments joining two nodes. To calculate the time-varying sensing constraints of the POI, a time-varying visibility volume (VV) captures airspace, sensing quality, and LOS constraints. As the POI moves and encounters new building geometry, its VV changes shape. To capture the variation of the VV, the road edges are sampled to form a space (and time) parameterized sequence of VVs. VV creation is described in our prior work [5, 6]. To balance the computational load of generating VVs and capturing changes in VV geometry, a bisection-inspired adaptive sampling method measures the similarity between two VVs, placing another VV in between if the geometry differs more than a specified rate. In this work, UAVs are approximated by a Dubins vehicle [7] operating at a constant altitude. The VVs are sliced at the UAV’s operating altitude, creating visibility polygons (VPs). Circles are inscribed in the VPs to produce static visibility orbits (VOs). Lastly, the time-varying VO is constructed by linearly interpolated neighboring static VOs. If the UAV remains on the time-varying VO, the UAV is inside the POI’s VV. To drive the UAV to converge to this time-varying VO a vector field provides a reference heading to a steering controller for the Dubins vehicle model.
2.1 Contributions
@@ 2.2 Paper Organization
The remainder of this paper is organized as follows. Section 3 discusses related work. Section 4 explains the environment model, vehicle motion model, point of interest motion, and measurement equations. Two problem statements that are the focus of this paper, reference orbit design and controller design, are given. Section 5 finds the feasibility of the visibility orbit for a Dubins vehicle and describes an algorithm for constructing the time-varying orbit inside the time-varying VPs. Another algorithm that adaptively places VVs along a road network to satisfy a desired maximum rate of change in geometry is also introduced. Section 6 derives the guidance vector field, and the steering controller that force the UAV onto the VO. Section 7 analyzes the numerical simulations that illustrate the proposed approach. Section 8 presents a flight test experiment. Section 9 concludes the paper.
3 Related Work
This section reviews work related to target tracking including environment and visibility models, control methodologies, and vehicle motion models. The section highlights recent work on target tracking in urban environments with occlusions. The relation of the proposed work to existing literature is discussed.
3.1 Environment and Visibility Models for Target Tracking
The POI’s environment determines how a UAV can position itself to view a POI. Many prior works consider environments that do not contain obstacles that may obscure a UAV’s visual sensors. For example, some authors assume that both the UAV and target move in a two-dimensional (2D) plane [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], the UAV moves in three-dimensions (3D) while the target is constrained to a horizontal plane [23, 24, 25, 26, 27, 28, 29, 30, 1], or both the UAV and target move in 3D [31, 32, 33]. Previous target tracking algorithms consider a target POI that moves over a road modeled either as a continuous region [34] or a discretized road network or group of cells [35, 36]. In [37], the UAV must avoid colliding with the urban environment, but the tracking method does not consider the obstruction of a visual sensor. In [36], the target POI is constrained to an urban road network, but occlusions are not considered. Researchers have also explored target tracking with more general non-building obstacles. In [38], a target moves in 2D and a team of multirotors moves in 3D while maneuvering to avoid obscuring another team member’s view of the target. Reference [33] considers spheres floating above the ground that UAVs must avoid and look beyond to track a target moving in 3D. The effect of mountainous terrain on target velocity is explored in [32]. Finally, some studies consider buildings and other obstacles blocking visual sensors prohibiting measurement of the target [35, 39, 40, 41, 42, 43]—the POI is blocked if there is an obstructed line from the camera to the target.
3.2 Vehicle Motion Models for Target Tracking
Multirotors and fixed-wing UAVs are commonly used in target-tracking applications. Multirotor UAVs can be modeled as single integrators [38, 11] or double integrators [37, 36, 44, 45]. More complex rigid-body dynamics have also been considered [24, 27, 28, 1] and are necessary when the vehicle’s attitude determines the onboard camera’s orientation. Fixed-wing UAVs are often modeled with turn rate constraints such as the acceleration limited double integrator [46, 18, 47], Dubins car [9, 35, 39, 10, 48, 41, 16, 30, 49, 21, 22], and 3D Dubins airplane [48, 41]. More sophisticated fixed-wing models are used when variable speed [34, 42, 50, 15] or roll angle [23] are relevant when an aircraft can change elevation to look past obstacles [43] or when the UAV is tracking another aircraft [31].
3.3 Target Tracking Control Methodologies
Methods for steering a vehicle to track a moving target can be grouped into (i) path-planning methods and (ii) feedback-control strategies. Path-planning methods explicitly plan a path over a fixed or moving time horizon to maintain the target in view. For example, [37, 35] investigated a single-step look ahead, and [8] investigated variable-step planners. The optimal path for a fixed-wing UAV performing standoff tracking can be found with dynamic programming [14] or using model predictive control (MPC). MPC-based methods have also been considered for multirotors [38, 33] and fixed-wing aircraft with various sensing constraints [17], fixed cameras [23, 49], and gimballed cameras [41, 42, 43]. In [10], a cycle that visits multiple targets is constructed with Dubins paths. In [1] $\rm A^{*}$ planning is used for tracking a target with a known path. Reference [13] uses the intersection of multiple visibility volumes to view multiple targets.
Feedback-control strategies define a control objective, such as maintaining a target-relative position or circulating a target, and employ error feedback in real time. One example is a Lyapunov vector field guidance law that steers fixed-wing UAVs onto fixed-radius-circular orbits around a target with constant velocity motion models [34, 15, 16, 30, 17], acceleration limited double integrator models [46, 18, 47], or single integrator models with a curvature constraint [19]. Lyapunov vector field control methods have also been adapted for a multirotor to track a target with a fixed camera [29, 51]. Other methods of forcing circulation on fixed-radius-circular orbits are sliding mode control—for example, driving a fixed-wing to circle a target [21]—and other nonlinear control laws that force circulation [52, 22]. Feedback control laws have also been proposed to steer fixed-wing UAVs [35, 17, 48, 10], and multirotors [28, 12, 36, 11] to positions relative to the target. In [44], sliding mode control forces a constant speed vehicle onto a pre-planned tracking trajectory. Neural-network-based approaches tracking targets by training a neural network with reinforcement learning [27, 53, 40] and approximating an optimal control set with a deep neural network [25].
3.4 Relation of Proposed Approach to Prior Work
This work extends our previous work in tour planning where a UAV images a series of static POIs in an urban environment [6]. Rather than multiple static view volumes, this work considers a single dynamic view volume. This work is similar to [43], where a cooperative multi-UAV system tracks a target through an urban environment. Similarly, this work uses a circulation guidance law to control the UAV onto a standoff trajectory as in [16]. However, our approach differs in the formulation of the control objective. Rather than ignoring the visibility of the target [16] or using model predictive control to account for visibility in [43], time-varying closed reference orbits are designed. The orbits are shown to be feasible for a Dubins vehicle and contained within the visibility volume, ensuring the POI is not occluded. A guidance vector field and a steering controller force the UAV onto the time-varying reference orbits. The guidance vector field is unique from previous work [54] because it is a constant magnitude vector field that accounts for both a dynamic orbit radius and a translating orbit. The steering controller is based on a novel Lyapunov function.
@@ 4 Problem Formulation
This section formulates the problem of tracking a moving target through an urban environment. The UAV’s motion is approximated by a Dubins vehicle, and its environment consists of 2.5D polygonal obstacles and a road network. The section introduces the POI’s motion on the road network, as well as the UAV’s sensor model and its measurement of the POI. Lastly, the problem statements are formulated as creating the time-varying VO and designing a controller to attract the UAV to the time-varying VO.
@@ 4.1 Vehicle Motion
where the system state is $\bm{q}\in{\rm SE}(2)$, the inertial planar position is $(q_{x},q_{y})$, the heading angle $q_{\psi}$ is measured counter-clockwise from the inertial positive $x$-axis, and the forward velocity $v$ is a constant value. The turn-rate control is bounded by
@@
where $u_{\psi,{\rm max}}$ is the maximum allowable turn rate. The bounded turn rate introduces a minimum turn radius $r_{\rm min}=v/u_{\psi,{\rm max}}$ and maximum curvature
$\kappa_{\rm max}=\frac{u_{\psi,{\rm max}}}{v}\;.$ (3)
The unsigned curvature of the trajectory of system (1) is
@@
The Dubins model is a good approximation for a fixed-wing aircraft [9, 35, 39, 10, 48, 41, 16, 30, 49, 21, 22] but can also model multirotors [55]. When used for multirotors, Dubins paths ensure smoother changes in velocity making them appropriate for high-speed path planning.

### v2

@@ Abstract
This paper considers the problem of tracking a point of interest (POI) moving along a known trajectory on the ground with an uncrewed aerial vehicle (UAV) modeled as a Dubins vehicle using a line-of-sight (LOS) sensor through an urban environment that may occlude the POI. A visibility volume (VV) encodes a time-varying, three-dimensional representation of the sensing constraints for a particular POI position. A constant-altitude, translating, and radially time-varying circular standoff orbit is then inscribed within the dynamically changing VV centered at the POI position. The time-varying VV is approximated by placing static VVs along the POI’s trajectory using an adaptive metric that restricts the volume change of consecutive VVs to below a specified rate. The time-varying circular standoff orbit is proven to be feasible for a Dubins vehicle and approximated with a piecewise set of linearly interpolated circular orbits inside the static VVs. A steering controller is derived that drives the UAV to the time-varying standoff orbit. Numerical simulations and a flight test illustrate the proposed approach.
@@ 1 Nomenclature
$(x,y,z)$  = inertial directions (m) easting, northing, and up
@@
$\psi$  = yaw angle (rad) about $z$ direction
$t$  = time (s)
@@
$v$  = forward velocity of the UAV (m/s)
$u_{\psi}$  = turn-rate control for the UAV (rad/s)
$\kappa$, $\kappa_{s}$  = signed and unsigned curvature of a trajectory (1/m)
$r_{\rm min}$  = minimum turn radius of the UAV (m); $1/\kappa_{\rm max}$
$h_{\rm UAV}$  = flight altitude of the UAV (m)
$B$  = set of $N_{B}$ 2.5-dimensional objects; $\{B_{0},\ldots,B_{N_{B}-1}\}$
$A_{i}$  = the order set of points representing the base of the $i$th object $B_{i}$
$h_{i}$  = the height of the $i$th object $B_{i}$ (m)
@@
$\bm{g}$  = inertial position of the moving point of interest (m); $[g_{x},g_{y}]^{\rm T}$
$v_{\rm g}$  = magnitude of the velocity (m/s) for the moving point of interest
$\gamma$  = direction (rad) of the point of interest’s velocity
@@
$d_{\rm max}$  = maximum viewing distance for imaging (m)
$\alpha(\bm{g})$, $\hat{\alpha}(\bm{g})$  = visibility volume and its approximation for target position $\bm{g}$
$h_{\rm building}$  = maximum height of the tallest building (m)
$h_{\rm feasible}$  = maximum height of the feasible airspace (m)
@@
$L(\tau;\bm{g},\bm{\rho})$  = equation of a line from $\bm{g}$ to $\bm{\rho}$ parameterized by $\tau$ in $\mathbb{R}^{3}$
$\theta$  = angle (rad) about the point $\bm{g}$ measured from a line parallel to the $x$ axis
@@
$R(t)$  = radius of a morphing visibility orbit (m)
$\bm{x}_{\mathcal{I}}$  = inertial trajectory (m) of the UAV in the $xy$ plane; $\bm{x}_{\mathcal{I}}\in\mathbb{R}^{2}$
$\bm{e}_{r}$,$\bm{e}_{\theta}$  = a polar coordinate system from the point $\bm{g}$
$r$  = UAV’s distance (m) from the POI
$d(\hat{\alpha}(\bm{g}_{i}),\hat{\alpha}(\bm{g}_{j}))$  = measure (m3) of the difference between two triangular meshes approximating visibility volumes
$d_{\rm cutoff}$  = maximum difference (m3) between two visibility volumes
$V_{1}$, $V_{2}$  = Lyapunov-like scalar functions
@@
$\bm{\xi}$  = inertial coordinates (m) for a kinematic particle; $\xi\in\mathbb{R}^{2}$
$\bm{u}$  = a vector field (m/s) that forces a particle onto an orbit $O(\theta,t)$; $\bm{u}\in\mathbb{R}^{2}$
$\bm{M}$  = a matrix component of the vector field $\bm{u}$
@@
$\psi_{d}$  = desired heading angle (rad)
@@ 2 Introduction
Uncrewed aerial vehicles (UAVs) are widely used for airborne sensing tasks that require maintaining a direct line of sight (LOS) to a target point of interest (POI). Applications of tracking a moving POI include cinematography, filming sporting events and parades, search and rescue, and police and military surveillance [1, 2]. For UAV tracking missions where a POI moves along a road through an urban environment, the directional sensor’s view of a POI may be obstructed by buildings, tunnels, bridges, trees, and other structures. This work creates a planning and control method to ensure a UAV maintains an unobstructed LOS to the a POI, moving along a known trajectory, while continuously orbiting it. By orbiting the POI, the UAV can observe it from various angles. Moreover, an encircling motion is required when the UAV cannot hover over the POI (e.g., if the POI’s speed is less than the UAV’s minimum speed). The POI moves over a road network represented as a graph with nodes representing points along roads and edges representing road segments joining two nodes. To calculate the time-varying sensing constraints of the POI, a time-varying visibility volume (VV) captures airspace, sensing quality, and LOS constraints. As the POI moves and encounters new building geometry, its VV changes shape. To capture the variation of the VV, the road edges are sampled to form a space (and time) parameterized sequence of VVs. VV creation is described in our prior work [3, 4]. To balance the computational load of generating VVs and capturing changes in VV geometry, a bisection-inspired adaptive discretization method measures the similarity between two VVs, placing another VV in between if the geometry differs more than a specified rate. In this work, a UAV is approximated by a Dubins vehicle [5] operating at a constant altitude. The VVs are sliced at the UAV’s operating altitude, creating visibility polygons (VPs). Circles are inscribed in the VPs to produce static visibility orbits (VOs). Lastly, the time-varying VO is constructed by linearly interpolating neighboring static VOs. If the UAV remains on the time-varying VO, the UAV is inside the POI’s VV. To drive the UAV to converge to this time-varying VO, a vector field provides a reference heading to a steering controller for the Dubins vehicle model.
The POI’s environment determines how a UAV can position itself to view a POI. Many prior works consider environments that do not contain obstacles that may obscure a UAV’s visual sensors. For example, some authors assume that both the UAV and target move in a two-dimensional (2D) plane [6, 7, 8], the UAV moves in three-dimensions (3D) while the target is constrained to a horizontal plane [1, 9], or both the UAV and target move in 3D [10, 11]. Previous target tracking algorithms consider a target POI that moves over a discretized road network or group of cells [12, 13, 14]. In [15], the UAV must avoid colliding with the urban environment, but the tracking method does not consider the obstruction of a visual sensor. In [13], the target POI is constrained to an urban road network, but occlusions are not considered. Researchers have also explored target tracking with more general non-building obstacles. In [16], a target moves in 2D and a team of multirotors moves in 3D while maneuvering to avoid obscuring another team member’s view of the target. Reference [11] considers spheres floating above the ground that UAVs must avoid and look beyond to track a target moving in 3D. The effect of mountainous terrain on target velocity is explored in [10]. Finally, some studies consider buildings and other obstacles blocking visual sensors prohibiting measurement of the target [12, 14, 17, 18]—the POI is blocked if there is an obstructed line from the camera to the target.
Multirotors and fixed-wing UAVs are commonly used in target-tracking applications. Multirotor UAVs can be modeled as single integrators [16] or double integrators [15, 13]. More complex rigid-body dynamics have also been considered [1] and are necessary when the vehicle’s attitude determines the onboard camera’s orientation. Fixed-wing UAVs are often modeled with turn rate constraints, such as the acceleration-limited double integrator [19], Dubins car [12, 14, 18, 7, 8, 9], and 3D Dubins airplane [18]. More sophisticated fixed-wing models are used when variable speed [20] or roll angle [21] are relevant when an aircraft can change elevation to look past obstacles [22].
Methods for steering a vehicle to track a moving target can be grouped into (i) path-planning methods and (ii) feedback-control strategies. Path-planning methods explicitly plan a path over a fixed or moving time horizon to maintain the target in view. For example, [15, 12] investigated a single-step look-ahead strategy. The optimal path for a fixed-wing UAV performing standoff tracking can be found with model predictive control (MPC). MPC-based methods have also been considered for multirotors [16, 11] and fixed-wing aircraft with fixed cameras [21], and gimballed cameras [18, 20, 22]. In [1] $\rm A^{*}$ planning is used for tracking a target with a known path. Feedback-control strategies define a control objective, maintaining a target-relative position or circulating a target, by employing error feedback. One example is a Lyapunov vector field guidance law that steers fixed-wing UAVs onto fixed-radius-circular orbits around a target with constant velocity motion models [9], acceleration limited double integrator models [19], or single integrator models with a curvature constraint [6]. Other methods of forcing circulation on fixed-radius-circular orbits are sliding mode control—for example, driving a fixed-wing to circle a target [7]—and other nonlinear control laws that force circulation [8]. Feedback control laws have also been proposed to steer fixed-wing UAVs [12], and multirotors [13] to positions relative to the target. Reinforcement learning is used to train a neural network to track a target in [17].
@@
The remainder of this paper is organized as follows. Section 3 explains the environment model, vehicle motion model, point of interest motion, and measurement model. The two problem statements in this work, reference orbit design and controller design, are given. Section 4 finds the feasibility of the visibility orbit for a Dubins vehicle and describes an algorithm to adaptively place VVs along a road network to satisfy a limit of geometry change. Another proposed algorithm constructs the time-varying orbit inside the time-varying VPs. Section 5 derives the guidance vector field and the steering controller that force the UAV onto the VO. Section 6 analyzes the numerical simulations illustrating the proposed approach. Section 7 presents a flight test experiment. Section 8 concludes the paper.
@@ 3 Problem Formulation
This section formulates the problem of tracking a moving target through an urban environment. The UAV’s motion is approximated by a Dubins vehicle, and its environment consists of 2.5D obstacles and a road network. The section introduces the POI’s motion, the UAV’s sensor model, and measurement of the POI. Lastly, the problem statements are formulated as creating the time-varying VO and designing a controller to attract the UAV to the time-varying VO.
@@ 3.1 Vehicle Motion
where the state is $\bm{q}\in{\rm SE}(2)$, the position is $(q_{x},q_{y})$, the heading angle $q_{\psi}$ is measured counter-clockwise from the inertial positive $x$-axis, and the forward velocity $v$ is a constant value. The turn-rate control is bounded by
@@
where $u_{\psi,{\rm max}}$ is the maximum allowable turn rate. The bounded turn rate creates a minimum turn radius $r_{\rm min}=v/u_{\psi,{\rm max}}$ and maximum curvature $\kappa_{\rm max}=u_{\psi,{\rm max}}/v$. The unsigned curvature of the trajectory of system (1) is
@@
The Dubins model can approximate fixed-wing aircraft [12, 14, 18, 9, 7, 8] and multirotors [23]. When used for multirotors, Dubins paths ensure smoother changes in velocity making them appropriate for high-speed path planning.